# On a B-field Transform of Generalized Complex Structures over Complex Tori

### v1

@@ 5.2 A deformation $DG_{X_{\mathcal{G}_{\tau}^{\nabla}}^{n}}$ of $DG_{X^{n}}$
on $U_{i}\cap U_{j}$ for each $i,j\in I$. Here, we take the trivial gerbe $\mathcal{G}$ with a non-trivial 0-connection $\{\nabla_{ij}=d+\omega_{ij}\}_{i,j\in I}$, and consider the deformation $M_{\mathcal{G}}$ of $M$ by this gerbe $\mathcal{G}$ (we only consider the case of the trivial gerbe with a non-trivial 0-connection since $\mathcal{G}_{\tau}^{\nabla}$ is trivial as a gerbe). Then, the family of transition functions $\{\varphi_{ij}\}_{i,j\in I}$ is preserved under this deformation since $\mathcal{G}$ is trivial, namely, the deformation $L^{\mathcal{G}}\to M_{\mathcal{G}}$ (as a bundle) of $L\to M$ by $\mathcal{G}$ is defined by the same family of transition functions $\{\varphi_{ij}^{\mathcal{G}}:=\varphi_{ij}\}_{i,j\in I}$. On the other hand, the twisted connection $\{\nabla_{i}^{\mathcal{G}}\}_{i\in I}$ on the deformed object $L^{\mathcal{G}}\to M_{\mathcal{G}}$ is defined by the relation
@@
$\tilde{\Omega}_{i}^{\mathcal{G}}:=d\omega_{i}^{\mathcal{G}}+\omega_{i}^{\mathcal{G}}\wedge\omega_{i}^{\mathcal{G}}=d\omega_{i}^{\mathcal{G}}.$
@@
on each $U_{i}$ by using the 1-connection $\{B_{i}\}_{i\in I}$ over $\mathcal{G}$ compatible with the 0-connection $\{\nabla_{i}^{\mathcal{G}}\}_{i\in I}$, there exists a unique 2-form $\Omega^{\mathcal{G}}$ on $M_{\mathcal{G}}$ with values in $\mathrm{End}(L^{\mathcal{G}})$ such that $\Omega^{\mathcal{G}}|_{U_{i}}=\Omega_{i}^{\mathcal{G}}$ for each $i\in I$. This 2-form $\Omega^{\mathcal{G}}$ is called the curvature form of the twisted connection $\{\nabla_{i}^{\mathcal{G}}\}_{i\in I}$ (see subsection 2.3 and subsection 2.4 in [30]).
Let us discuss the deformation of smooth complex line bundles with connections $E_{(s,a,q)}^{\nabla}\to X^{n}$, namely, we define the twisted smooth complex line bundle $E_{(s,a,q)}^{\tau}\to X_{\mathcal{G}_{\tau}^{\nabla}}^{n}$ with the twisted connection $\nabla_{(s,a,q)}^{\tau}$ as the deformation of a given smooth complex line bundle $E_{(s,a,q)}^{\nabla}\to X^{n}$ by $\mathcal{G}_{\tau}^{\nabla}$. The transition functions of $E_{(s,a,q)}^{\tau}$ coincide with the transition functions of $E_{(s,a,q)}$ since $\mathcal{G}(I,\mathcal{O},\xi)$ is trivial. On the other hand, the connection $\nabla_{(s,a,q)}$ on $E_{(s,a,q)}$ is locally deformed to the following twisted connection $\nabla_{(s,a,q)}^{\tau}$ (see also the definition (17)):
@@
In fact, we can easily check that this $\nabla_{(s,a,q)}^{\tau}$ is compatible with the transition functions of $E_{(s,a,q)}^{\tau}$, i.e., the relations of the form (20) hold. Similarly as in the case of $E_{(s,a,q)}^{\nabla}$, for simplicity, we sometimes denote this deformed object by
@@
from the 1-form (21) by considering its derivative, and this 2-form (22) is the 1-connection over $\mathcal{G}(I,\mathcal{O},\xi)$ itself compatible with the 0-connection $\nabla_{\tau}$ (see the local expression (14) and the definition (15)). Sometimes the 2-form
@@
in the local expression (22) is also called the B-field as mentioned in section 4. Here, we discuss the holomorphicity of $E_{(s,a,q)}^{\nabla}(\tau)$. We see that the following proposition holds.
@@
in the local expression (23) is the 1-connection over $\mathcal{G}(I,\mathcal{O},\xi)$ compatible with the 0-connection $\nabla_{\tau}$ (B-field). In general, for a given twisted smooth complex vector bundle with a twisted connection, such a twisted connection is an twisted integrable connection on the given twisted smooth complex vector bundle if and only if the (0,2)-part of its curvature form vanishes (see Lemma 2.18 and Lemma 2.19 in [30]). By direct calculations, the (0,2)-part of $\Omega_{(s,a,q)}^{\tau}$ turns out that
@@
$\mathrm{Hom}_{DG_{X_{\mathcal{G}_{\tau}^{\nabla}}^{n}}}\left(E_{(s^{1},a^{1},q^{1})}^{\nabla}(\tau),E_{(s^{2},a^{2},q^{2})}^{\nabla}(\tau)\right):=\bigoplus_{r\in\mathbb{N}\cup\{0\}}\mathrm{Hom}_{DG_{X_{\mathcal{G}_{\tau}^{\nabla}}^{n}}}^{r}\left(E_{(s^{1},a^{1},q^{1})}^{\nabla}(\tau),E_{(s^{2},a^{2},q^{2})}^{\nabla}(\tau)\right),$
$\mathrm{Hom}_{DG_{X_{\mathcal{G}_{\tau}^{\nabla}}^{n}}}^{r}\left(E_{(s^{1},a^{1},q^{1})}^{\nabla}(\tau),E_{(s^{2},a^{2},q^{2})}^{\nabla}(\tau)\right):=\Gamma\left(E_{(s^{1},a^{1},q^{1})}^{\nabla}(\tau),E_{(s^{2},a^{2},q^{2})}^{\nabla}(\tau)\right)\bigotimes_{C^{\infty}\left(X_{\mathcal{G}_{\tau}^{\nabla}}^{n}\right)}\Omega^{(0,r)}\left(X_{\mathcal{G}_{\tau}^{\nabla}}^{n}\right),$
@@
$d_{12}^{r}\left(\left\{\phi_{(l;m)}^{12}\right\}_{(l;m)\in I}\right):=\left\{\left(\nabla_{(s^{2},a^{2},q^{2})}^{\tau}\right)^{(0,1)}\phi_{(l;m)}^{12}-(-1)^{r}\phi_{(l;m)}^{12}\left(\nabla_{(s^{1},a^{1},q^{1})}^{\tau}\right)^{(0,1)}\right\}_{(l;m)\in I}$
@@
for each $(r_{1},r_{2})\in(\mathbb{N}\cup\{0\})^{2}$ by
$m\left(\left\{\phi_{(l;m)}^{23}\right\}_{(l;m)\in I},\left\{\phi_{(l;m)}^{12}\right\}_{(l;m)\in I}\right):=\left\{\phi_{(l;m)}^{23}\wedge\phi_{(l;m)}^{12}\right\}_{(l;m)\in I}.$
@@
As a corollary of Theorem 5.5, we see that holomorphic line bundles with integrable connections $E_{(s,a,q)}^{\nabla}$ on $X^{n}$ forms a dg-category $DG_{X^{n}}:=DG_{X_{\mathcal{G}_{O}^{\nabla}}^{n}}$ when we consider the case that $\mathcal{G}_{\tau}^{\nabla}$ is trivial, i.e., $\tau=O$, and such a dg-category is actually used in various studies (see [16, 22] etc.). In particular, although we can regard the dg-category $DG_{X_{\mathcal{G}_{\tau}^{\nabla}}^{n}}$ as the deformation of the dg-category $DG_{X^{n}}$ associated to the deformation from $X^{n}$ to $X_{\mathcal{G}_{\tau}^{\nabla}}^{n}$ by $\mathcal{G}_{\tau}^{\nabla}$, this can also be understood as the equivalence
@@
These facts are explained in [6] for example (cf. [17, 18]). In this paper, although we do not assume that $X^{n}$ is an abelian variety, namely, we regard $X^{n}$ as a general complex torus, it seems that the deformation (25) can also be interpreted as an analogue of the equivalence (26).

### v2

@@ 5.2 A deformation $DG_{X_{\mathcal{G}_{\tau}^{\nabla}}^{n}}$ of $DG_{X^{n}}$
on $U_{i}\cap U_{j}$ for each $i,j\in I$. Here, we take the trivial gerbe $\mathcal{G}$ with a non-trivial 0-connection $\{\nabla_{ij}=d+\omega_{ij}\}_{i,j\in I}$, and consider the deformation $M_{\mathcal{G}}$ of $M$ by this gerbe $\mathcal{G}$ (we only consider the case of the trivial gerbe with a non-trivial 0-connection since $\mathcal{G}_{\tau}^{\nabla}$ is trivial as a gerbe). Then, the family of transition functions $\{\varphi_{ij}\}_{i,j\in I}$ is preserved under this deformation since $\mathcal{G}$ is trivial. This result indicates that the deformation $L^{\mathcal{G}}\to M_{\mathcal{G}}$ (as a bundle) of $L\to M$ by $\mathcal{G}$ is defined by the same family of transition functions $\{\varphi_{ij}^{\mathcal{G}}:=\varphi_{ij}\}_{i,j\in I}$. In other words, $L^{\mathcal{G}}$ is an $id$-twisted smooth complex line bundle on $M_{\mathcal{G}}$ in the sense of [31, Definition 2.1] [^6] Although the twisting of $L$ by $\mathcal{G}$ is actually trivial since $\mathcal{G}$ is trivial, we regard the deformation $L^{\mathcal{G}}$ of $L$ as an $id$-twisted line bundle in order to specify that the deformation of $L$ is caused by the flat gerbe $(\mathcal{G},\{\nabla_{ij}\}_{i,j\in I})$.. On the other hand, the twisted connection $\{\nabla_{i}^{\mathcal{G}}\}_{i\in I}$ on the deformed object $L^{\mathcal{G}}\to M_{\mathcal{G}}$ is defined by the relation
@@
$\tilde{\Omega}_{i}^{\mathcal{G}}=d\omega_{i}^{\mathcal{G}}+\omega_{i}^{\mathcal{G}}\wedge\omega_{i}^{\mathcal{G}}=d\omega_{i}^{\mathcal{G}}.$
@@
on each $U_{i}$ by using the 1-connection $\{B_{i}\}_{i\in I}$ over $\mathcal{G}$ which is compatible with the 0-connection $\{\nabla_{i}^{\mathcal{G}}\}_{i\in I}$, there exists a unique 2-form $\Omega^{\mathcal{G}}$ on $M_{\mathcal{G}}$ with values in $\mathrm{End}(L^{\mathcal{G}})$ such that $\Omega^{\mathcal{G}}|_{U_{i}}=\Omega_{i}^{\mathcal{G}}$ for each $i\in I$. This 2-form $\Omega^{\mathcal{G}}$ is called the curvature form of the twisted connection $\{\nabla_{i}^{\mathcal{G}}\}_{i\in I}$ (see subsection 2.3 and subsection 2.4 in [31]).
Let us discuss the deformation of smooth complex line bundles with connections $E_{(s,a,q)}^{\nabla}\to X^{n}$, namely, we define the $id$-twisted smooth complex line bundle $E_{(s,a,q)}^{\tau}\to X_{\mathcal{G}_{\tau}^{\nabla}}^{n}$ with the twisted connection $\nabla_{(s,a,q)}^{\tau}$ as the deformation of a given smooth complex line bundle with a connection $E_{(s,a,q)}^{\nabla}\to X^{n}$ by $\mathcal{G}_{\tau}^{\nabla}$. The transition functions of $E_{(s,a,q)}^{\tau}$ coincide with the transition functions of $E_{(s,a,q)}$ since $\mathcal{G}(I,\mathcal{O},\xi)$ is trivial. On the other hand, the connection $\nabla_{(s,a,q)}$ on $E_{(s,a,q)}$ is locally deformed to the following twisted connection $\nabla_{(s,a,q)}^{\tau}$ (see also the definition (17)):
@@
In fact, we can easily check that this $\nabla_{(s,a,q)}^{\tau}$ is compatible with the transition functions of $E_{(s,a,q)}^{\tau}$, i.e., the relations of the form (19) hold. Similarly as in the case of $E_{(s,a,q)}^{\nabla}$, we sometimes denote this deformed object by
@@
from the 1-form (20) by considering its derivative, and this 2-form (21) is the 1-connection over $\mathcal{G}(I,\mathcal{O},\xi)$ itself which is compatible with the 0-connection $\nabla_{\tau}$ (see the local expression (14) and the definition (15)). The 2-form
@@
in the local expression (21) is also called a B-field as mentioned in section 4. Here, we discuss the holomorphicity of $E_{(s,a,q)}^{\nabla}(\tau)$. We see that the following proposition holds.
@@
$\frac{\partial s}{\partial x}(x)T=\left(\frac{\partial s}{\partial x}(x)T\right)^{t}$ (22)
@@
in the local expression (23) is the 1-connection over $\mathcal{G}(I,\mathcal{O},\xi)$ which is compatible with the 0-connection $\nabla_{\tau}$ (a B-field). As mentioned in Remark 5.2, for an $\alpha$-twisted smooth complex vector bundle with a twisted connection (this $\alpha$ denotes the 2-cocycle which defines a given gerbe), we call the twisted connection a twisted integrable connection on the $\alpha$-twisted smooth complex vector bundle if the (0,2)-part of its curvature form vanishes (see [31, Lemma 2.18, Lemma 2.19]). By direct calculations, the (0,2)-part of $\Omega_{(s,a,q)}^{\tau}$ turns out that
@@
$\mathrm{Hom}_{DG_{X_{\mathcal{G}_{\tau}^{\nabla}}^{n}}}\left(E_{(s^{1},a^{1},q^{1})}^{\nabla}(\tau),E_{(s^{2},a^{2},q^{2})}^{\nabla}(\tau)\right)=\bigoplus_{r\in\mathbb{N}\cup\{0\}}\mathrm{Hom}_{DG_{X_{\mathcal{G}_{\tau}^{\nabla}}^{n}}}^{r}\left(E_{(s^{1},a^{1},q^{1})}^{\nabla}(\tau),E_{(s^{2},a^{2},q^{2})}^{\nabla}(\tau)\right),$
$\mathrm{Hom}_{DG_{X_{\mathcal{G}_{\tau}^{\nabla}}^{n}}}^{r}\left(E_{(s^{1},a^{1},q^{1})}^{\nabla}(\tau),E_{(s^{2},a^{2},q^{2})}^{\nabla}(\tau)\right)=\Gamma\left(E_{(s^{1},a^{1},q^{1})}^{\nabla}(\tau),E_{(s^{2},a^{2},q^{2})}^{\nabla}(\tau)\right)\bigotimes_{C^{\infty}\left(X_{\mathcal{G}_{\tau}^{\nabla}}^{n}\right)}\Omega^{(0,r)}\left(X_{\mathcal{G}_{\tau}^{\nabla}}^{n}\right),$
@@
$d_{12}^{r}\left(\left\{\phi_{(l;m)}^{12}\right\}_{(l;m)\in I}\right)=\left\{\left(\nabla_{(s^{2},a^{2},q^{2})}^{\tau}\right)^{(0,1)}\phi_{(l;m)}^{12}-(-1)^{r}\phi_{(l;m)}^{12}\left(\nabla_{(s^{1},a^{1},q^{1})}^{\tau}\right)^{(0,1)}\right\}_{(l;m)\in I}$
@@
for each $(r_{1},r_{2})\in(\mathbb{N}\cup\{0\})\times(\mathbb{N}\cup\{0\})$ by
$m\left(\left\{\phi_{(l;m)}^{23}\right\}_{(l;m)\in I},\left\{\phi_{(l;m)}^{12}\right\}_{(l;m)\in I}\right)=\left\{\phi_{(l;m)}^{23}\wedge\phi_{(l;m)}^{12}\right\}_{(l;m)\in I}.$
@@
As a corollary of Theorem 5.7, we see that holomorphic line bundles with integrable connections $E_{(s,a,q)}^{\nabla}$ on $X^{n}$ forms a dg-category $DG_{X^{n}}:=DG_{X_{\mathcal{G}_{O}^{\nabla}}^{n}}$ when we consider the case that $\mathcal{G}_{\tau}^{\nabla}$ is trivial, i.e., $\tau=O$, and actually, such a dg-category is used in various studies (see [17, 23] etc.). We can regard the dg-category $DG_{X_{\mathcal{G}_{\tau}^{\nabla}}^{n}}$ as the deformation of the dg-category $DG_{X^{n}}$ associated to the deformation from $X^{n}$ to $X_{\mathcal{G}_{\tau}^{\nabla}}^{n}$ by $\mathcal{G}_{\tau}^{\nabla}$, this can also be understood as the equivalence
@@
These facts are explained in [7] for example (cf. [18, 19]). In this paper, although we do not assume that $X^{n}$ is an abelian variety (we regard $X^{n}$ as a general complex torus), it seems that the deformation (25) can also be interpreted as an analogue of the equivalence (26).
5.3 A problem
The goal of subsection 5.3 and subsection 5.4 is to construct the deformation of objects $\mathscr{L}_{(s,a,q)}^{\nabla}\in Fuk_{\rm sub}(\check{X}^{n})$ associated to the deformation from $\check{X}^{n}$ to $\check{X}_{\mathcal{G}_{\tau}^{\nabla}}^{n}$. In particular, we need to consider the deformation of unitary local systems $\mathcal{L}_{(s,a,q)}^{\nabla}$ in order to achieve this goal. However, when we discuss the deformation of objects $\mathscr{L}_{(s,a,q)}^{\nabla}\in Fuk_{\rm sub}(\check{X}^{n})$, in general, there exists an alternating matrix $\tau\in M(n;\mathbb{R})$ such that any deformations of $\mathcal{L}_{(s,a,q)}^{\nabla}$ cannot be constructed in the framework of the usual Fukaya categories. The purposes of this subsection are to explain this problem, and to give the outline of an idea to solve it.
Let us again recall the usual definition of objects of the Fukaya category over a given symplectic manifold $M$ with a B-field $2\pi B$. An object of the Fukaya category over this complexified symplectic manifold consists of a Lagrangian submanifold $L$ in $M$ and a smooth complex line bundle $\mathcal{L}\to L$ with a unitary connection $\nabla_{\mathcal{L}}$ such that
$\Omega_{\nabla_{\mathcal{L}}}=\Bigl{.}-2\pi\mathbf{i}B\Bigr{|}_{L},$ (27)
where $\Omega_{\nabla_{\mathcal{L}}}$ denotes the curvature form of $\nabla_{\mathcal{L}}$. In particular, there exists such a unitary connection $\nabla_{\mathcal{L}}$ if and only if $[B]\in H^{2}(L;\mathbb{Z})$, namely, there does not exist a unitary connection $\nabla_{\mathcal{L}}$ which satisfies the condition (27) on each Lagrangian submanifold $L$ such that $[B]\not\in H^{2}(L;\mathbb{Z})$.
For later convenience, we prepare some notations. Note that $\tau\in M(n;\mathbb{R})$ is an alternating matrix. For each component $\tau_{ij}$ of $\tau\in M(n;\mathbb{R})$ such that $1\leq i<j\leq n$, we set
$\lfloor\tau_{ij}\rfloor:=\mathrm{max}\Bigl{\{}m\in\mathbb{Z}\ |\ m\leq\tau_{ij}\Bigr{\}},$
and let us define two alternating matrices $\tau_{\mathbb{Z}}\in M(n;\mathbb{Z})$ and $\tau_{\rm frac}\in M(n;\mathbb{R})$ by
$(\tau_{\mathbb{Z}})_{ij}=\lfloor\tau_{ij}\rfloor,\ (\tau_{\mathbb{Z}})_{ji}=-\lfloor\tau_{ij}\rfloor,$
$(\tau_{\rm frac})_{ij}=\tau_{ij}-\lfloor\tau_{ij}\rfloor,\ (\tau_{\rm frac})_{ji}=-\tau_{ij}+\lfloor\tau_{ij}\rfloor,$
respectively. Then, we have
$\tau=\tau_{\mathbb{Z}}+\tau_{\rm frac}.$
By using these notations, the B-field $B_{\tau}^{\vee}$ on $\check{X}_{\mathcal{G}_{\tau}^{\nabla}}^{n}$ is expressed locally as
$B_{\tau}^{\vee}=B^{\vee}+\pi d\check{x}^{t}\tau d\check{x}=B^{\vee}+\pi d\check{x}^{t}\tau_{\mathbb{Z}}d\check{x}+\pi d\check{x}^{t}\tau_{\rm frac}d\check{x}.$
Here, we consider the case $\tau\in M(n;\mathbb{Z})$, i.e., $\tau_{\rm frac}=O$. Then, for the B-field $B^{\vee}+\pi d\check{x}^{t}\tau_{\mathbb{Z}}d\check{x}$, since
.$\Bigl{.}B^{\vee}\Bigr{|}_{L_{(s,a)}}=2\pi d\check{x}^{t}B_{\rm mat}^{\vee}\frac{\partial s}{\partial\check{x}}(\check{x})d\check{x},$
if we take the Lagrangian submanifold $L_{(s,a)}$ in $(\mathbb{R}^{2n}/\mathbb{Z}^{2n},\omega^{\vee})$ which is determined by each $s$ satisfying the relation
$B_{\rm mat}^{\vee}\frac{\partial s}{\partial\check{x}}(\check{x})=\left(B_{\rm mat}^{\vee}\frac{\partial s}{\partial\check{x}}(\check{x})\right)^{t},$
then the requirement (27) is satisfied: [^8] It is natural to consider that $L_{(s,a)}$ is not deformed since $\omega^{\vee}$ is preserved in the deformation from $\check{X}^{n}$ to $\check{X}_{\mathcal{G}_{\tau}^{\nabla}}^{n}$.
$\left[d\check{x}^{t}B_{\rm mat}^{\vee}d\check{y}+\frac{1}{2}d\check{x}^{t}\tau_{\mathbb{Z}}d\check{x}\right]=\left[\frac{1}{2}d\check{x}^{t}\tau_{\mathbb{Z}}d\check{x}\right]\in H^{2}(L_{(s,a)};\mathbb{Z}).$
Hence, when we focus on such Lagrangian submanifolds $L_{(s,a)}$, we can expect that the deformation of objects $\mathscr{L}_{(s,a,q)}^{\nabla}\in Fuk_{\rm sub}(\check{X}^{n})$ can be constructed in the framework of the usual Fukaya categories.
However, in the case $\tau\not\in M(n;\mathbb{Z})$, i.e., $\tau_{\rm frac}\not=O$, for the B-field $B^{\vee}+\pi d\check{x}^{t}\tau d\check{x}$, it is easy to check that
$\left[d\check{x}^{t}B_{\rm mat}^{\vee}d\check{y}+\frac{1}{2}d\check{x}^{t}\tau d\check{x}\right]\not\in H^{2}(L_{(s,a)};\mathbb{Z})$
holds for any Lagrangian submanifold $L_{(s,a)}$ in $(\mathbb{R}^{2n}/\mathbb{Z}^{2n},\omega^{\vee})$ even though we can consider the deformation $E_{(s,a,q)}^{\nabla}(\tau)$ of each $E_{(s,a,q)}^{\nabla}$ depending on $\tau\not\in M(n;\mathbb{Z})$. Thus, we can not discuss the deformation of objects $\mathscr{L}_{(s,a,q)}^{\nabla}\in Fuk_{\rm sub}(\check{X}^{n})$ in the usual framework over $\check{X}_{\mathcal{G}_{\tau}^{\nabla}}^{n}$ under the assumption $\tau\not\in M(n;\mathbb{Z})$.
Our proposal to solve this problem is to consider the twisting of the complexified symplectic torus $(\mathbb{R}^{2n}/\mathbb{Z}^{2n},\omega^{\vee}-\mathbf{i}(B^{\vee}+\pi d\check{x}^{t}\tau_{\mathbb{Z}}d\check{x}))$ by the flat gerbe $\check{\mathcal{G}}_{\tau_{\rm frac}}^{\nabla}$ associated to $\tau_{\rm frac}\in M(n;\mathbb{R})$. In other words, our idea is to employ “twisted” line bundles associated to the flat gerbe $\check{\mathcal{G}}_{\tau_{\rm frac}}^{\nabla}$ instead of “usual” line bundles on Lagrangian submanifolds. Strictly speaking, we do not discuss the compatibility of our proposal and some structures required to define (an analogue of) the Fukaya category over $\check{X}_{\check{\mathcal{G}}_{\tau}^{\nabla}}^{n}$, so in this sense, our proposal is just an idea.